OpenStudy (anonymous):

dy/dx=(y+5)(x+2) and y=1 when x=0 ?

OpenStudy (p0sitr0n):

dy/(y+5)=dx*(x+2) ln(y+5)=1/2(x+2)^2 y+5=e^(1/2*x+2)^2) y=e^(1/2*x+2)^2)-5